OpenStudy (anonymous):

find the domain in interval notation of a) f(x)= 5x-4/3x-2x b) f(x)= sqr 2x+7 c)3x^2-7x+5=11

jhonyy9 (jhonyy9):

a. on denominat is correct ???

jhonyy9 (jhonyy9):

b. all real numbers

jhonyy9 (jhonyy9):

a. sorry want to be denominator

OpenStudy (anonymous):

yes the denominator

jhonyy9 (jhonyy9):

c. 3x2 -7x+5=11 3x2 -7x -6 =0 x_1,_2 =(7+/- sqrt(49+72))/6 = (7+/- sqrt(121))/6 =(7+/- 11)/6 = - 2/3 and 3 (x+(2/3))(x-3)=0

jhonyy9 (jhonyy9):

a. (5x-4)/x so rstly we need suppose that x not can being 0 because than the fraction is undefined so than x can be all real numbers without zero

OpenStudy (anonymous):

so 5x-4/3-2x is undefined